OpenStudy (anonymous):

Greta had 5/8 pound of ham. She used 1/3 pound for sandwiches. How many pounds of ham dose she have left? A. 5/24 lb B. 7/24 lb C. 1/3 lb D. 4/5 lb

OpenStudy (anonymous):

@Mystiz

jabez177 (jabez177):

\[\frac{ 5\times 3 }{ 8\times 3 }\]

jabez177 (jabez177):

Whatever you do to the bottom, you do to the top. What is 5 x 3 and what is 8 x 3?

OpenStudy (anonymous):

5x3 is 15 and 8x3 is 24

jabez177 (jabez177):

Good now what x 3 will get you 24?

jabez177 (jabez177):

Yes, now what you do to the bottom you have to do to the top. 1 x 8 = ? - 3 x 8 = ?

OpenStudy (anonymous):

8/24?

OpenStudy (anonymous):

It is C. right

jabez177 (jabez177):

Yes, now the new fractions we have are 15 8 - and - 24 24

jabez177 (jabez177):

Now the question asks, "How much does she have left?" So is it an Addition question or Subtraction?

OpenStudy (anonymous):

Subtraction so 7/24 so it is B.

jabez177 (jabez177):

Great job! :D Do you think you can do it by yourself now?

OpenStudy (anonymous):

Yes Thank you

jabez177 (jabez177):

You are most welcome! :D